Find the equation of the straight line passing through (1, 2, 3) and perpendicular to the plane x + 2y – 5z + 9 = 0 ?

Grade:12

1 Answers

SHAIK AASIF AHAMED
9 years ago
Hello student,
Please find the answer to your question below
The equation of the straight line passing through (1, 2, 3) and perpendicular to the plane x + 2y – 5z + 9 = 0 is given by
1(x-1)+2(y-2)-5(z-3)=0
x+2y-5z+10=0